how do i solve using one of the 6 sum/diff identities and not the calculator?
tan(5pi/3 - pi/4)
: how do i solve using one of the 6 sum/diff identities and not the calculator?
tan(5pi/3 - pi/4)

Answer by stanbon(19007)

tan(A−B) = (tan A − tan B) / (1 + tan A tan B)

--------------------------------------------------------------------------------
tan(5pi/3 - pi/4) = (tan(5pi/3) - tan(pi/4)) / (1 + (tan(5pi/3)(tan(pi/4))
= (tan(-pi/3) - 1) / (1 + (tan(-pi/3)*1)
= (-sqrt(3)-1) / (1 + (-sqrt(3)*1)
= -(1+sqrt(3))/(1 - sqrt(3))
= [-(1+sqrt(3))^2]/ (-2)
= 3.7321...
=============================
Cheers,
Stan H.